### v2

# Approximate solutions to a nonlinear functional differential equation

Abstract
A functional differential equation related to the logistic equation is studied by a combination of numerical and perturbation methods. Parameter regions are identified where the solution to the nonlinear problem is approximated well by known series solutions of the linear version of the equation. The solution space for a certain class of functions is then mapped out using a continuation approach.
keywords:
Functional differential equation, bifurcation, Laguerre spectral collocation
{MSCcodes}
34K18, 34K28, 74S25
1 Introduction
The functional differential equation
(1) $u^{\prime}(t)+a\,u(t)=b\,u(\alpha t),\quad t>0,$
with $a$, $b$, and $\alpha>0$ constants, was the topic of two major papers, [10] and [16], both dating back about half a century. An application mentioned in the former of these references is the dynamics of an overhead current collection system for an electric locomotive. Even earlier papers derived the equation as a probabilistic model of absorption processes [1, 12]. In later years the equation has re-appeared in mathematical biology, such as in the modeling of cell growth [15, 19].
In this paper we present results for a nonlinear version of this equation,
(2) $u^{\prime}(t)+u(t)=u^{2}(\alpha t),\quad t>0,$
considered by [2] and more recently by [7]. The latter paper, in which the term $\alpha$-Riccati equation is coined, discusses an application to cascading processes in Fourier space representations of some nonlinear PDEs (including the Navier–Stokes equations). The two references cited both approach the nonlinear problem from a probabilistic viewpoint, in particular a stochastic branching process. Here, the tool is numerical computation supported by a perturbation analysis.
For $\alpha=1$, both (1) and (2) reduce to well-known equations that will not be considered here. For $\alpha<1$, the equations are delay differential equations with proportional or ‘pantograph-type’ delay. This regime was investigated in [10] and [16] in the linear case, and [7] in the nonlinear case. In particular, existence and uniqueness of solutions were established for both (1) and (2).
The case of $\alpha>1$ in (1) and (2) involves a ‘time-advanced’ argument, however, in applications this typically represents a non-temporal variable, thus avoiding a violation of causality in the usual time-evolution sense. For example, the independent variable may represent a scale in the biological settings treated in [15], the magnitude of the wave vector in the self-similar Navier–Stokes equations as discussed in [7], or a transparency scale as in [1]. In the pantograph-type systems introduced in [10], and carefully analyzed in [16], the parameter $\alpha$ is also allowed to take values greater than unity for modeling purposes. Further examples in which $\alpha>1$ arise are models of nerve axons from conduction theory [4] and preview control from control theory, where it models anticipation [3]. As the simplest nontrivial nonlinear equation with this feature, we investigate solutions to (2) in order to provide insight into more complex time-advanced functional differential equations arising in applications.
While [10] and [16] give a thorough analysis of properties of solutions of the linear equation (1), current knowledge for the non-linear counterpart (2) is incomplete. A visual status report is given in [7, Fig. 2]. This diagram catalogs the existence and uniqueness properties of solutions in several regions of the $(\alpha,u_{0})$ plane, where $u(0)=u_{0}$ represents an initial condition. In several regions either existence or uniqueness remains unknown. The focus in this manuscript will be on solutions to (2) for the case $\alpha>1,$ $u_{0}=1$, which forms a small but important section of that diagram.
We shall make no attempt to prove any theorems regarding existence/uniqueness, but rather focus on quantitative aspects such as the approximation of solutions, or, since multiple solutions exist, to determine how many solutions in a certain class are possible for each value of $\alpha$. By a combination of analysis and computation we shall identify parameter regions where the solutions to the nonlinear problem (2) can be approximated well by certain explicitly known series solutions of the linear problem (1).
For numerical solutions we use a spectral collocation method for functional differential equations. (Details can be found in [14] and a short description in section 4.) This method provides more efficient and more accurate results than other methods reported in the literature. Two such methods for the linear problem (1) are described in [10]: A finite difference method, applicable to the case $\alpha<1$, and a Lanczos tau-method for both $\alpha<1$ and $\alpha>1$. Numerical results, of low accuracy in today’s terms but impressive for 1970s technology, are given. The only numerical approach to the nonlinear problem (2) that we are aware of is the probabilistic simulation of [5]. Like most methods of Monte Carlo-type, however, it is not efficient for doing the extensive experimentation that will be reported here. Because the problem (2) is nonlinear, any numerical procedure would require iteration and this demands good starting guesses. Such approximations, which may be of interest in their own right, are obtained here by an elementary perturbation analysis.
The outline of this paper is therefore as follows. In section 2 we introduce some aspects of the linear problem (1) that are pertinent in the analysis of the nonlinear problem (2). In particular, we consider $a=1$ and $b=2$, corresponding to the linearisation of (2) about the constant solution $u=1$, and demonstrate the existence of nontrivial solutions to the homogeneous linear problem for certain characteristic values of $\alpha$, designated here as $\alpha_{n}$, $n=1,2,\ldots$. In section 3 we show that these characteristic solutions, appropriately scaled, provide the leading order term in a perturbation of the constant solution to (2) in the neighbourhood of these characteristic values. In section 4 we use these perturbation approximations in combination with the spectral collocation method and pseudo-arclength continuation to determine numerically a family of solutions to (2) for a range of $\alpha$ values, which we discuss in section 5.
2 Summary of results for the linear equation
The papers [10] and [16] contain a wealth of information on the linear equation (1). This section highlights only the major results that will be used in the nonlinear analysis of the next section. In particular, in that section the values $a=1$ and $b=2$ will be relevant, i.e.,
(3) $u^{\prime}(t)+u(t)=2u(\alpha t).$
This equation has the series solution
(4) $u(t)=CE(t;\alpha),\quad\text{where}\quad E(t;\alpha)=e^{-t}\Big{(}1+\sum_{k=1}^{\infty}\frac{2^{k}e^{(1-\alpha^{k})t}}{(1-\alpha)(1-\alpha^{2})\cdots(1-\alpha^{k})}\Big{)},$
and $C$ is a constant. This solution is a special case of one given in [16] for (1). [^1] Neither [16] nor [10] explicitly derive (4) but [15] and [19] contain derivations based on Laplace transforms and Mellin transforms, respectively. The series converges absolutely and uniformly for all $t\geq 0$ in the case $\alpha>1$.
The initial condition $u(0)=u_{0}$ can be matched to (4) by setting $t=0$, but of particular interest is the case $u_{0}=0$. This can be satisfied by setting $C=0$, but nontrivial solutions are possible for special values of $\alpha$, as the following lemma shows. These values of $\alpha$ can therefore be viewed as eigenvalues or characteristic values of the linear operator. In other words, for these values of $\alpha$ the function (4) lies in the nullspace of (3).
Lemma 1: The equation
(5) $01+\sum_{k=1}^{\infty}\frac{2^{k}}{(1-\alpha)(1-\alpha^{2})\cdots(1-\alpha^{k})}=$ 0
is satisfied if and only if $\alpha=\alpha_{n}=2^{1/n}$ for some positive integer $n$.
Proof: An identity well-known in the field of $q$-series [11, p. 11] states that
(6) $11+\sum_{k=1}^{\infty}\frac{(-1)^{k}q^{k(k+1)/2}c^{k}}{(1-q)(1-q^{2})\cdots(1-q^{k})}=(1-cq)(1-cq^{2})\cdots(1-cq^{n})\cdots,\quad|q|<$ 1 .
The proof is completed by substituting $c=2$, $q=1/\alpha$.
A related observation was made in [16], namely that the initial condition $u(0)=0$ “can occur only for exceptional pairs $a$ and $b$” (with these constants as defined in (1)). The authors did not explore further, but in [19, Eq. (2.13)] it was shown that $u(0)=0$ if and only if $b=a\alpha^{n}$ for some integer $n$. This follows from (6) by substituting $c=b/a$ and $q=1/\alpha$. Note that both [16] and [19] addressed the situation with $\alpha$ fixed; our interest is in the case when $a$ and $b$ are fixed.
To distinguish these cases we shall call the functions (4) with $\alpha=\alpha_{n}=2^{1/n}$ “characteristic functions” rather than “eigenfunctions” as in [19]. The first few characteristic functions of (3) are shown in Fig. 1, normalised such that $\|u\|_{2}=1$ (the norm is the standard, unweighted $L^{2}$ norm on $[0,\infty)$.)
Figure 1: $L^{2}$ normalised characteristic solutions to (3), as defined by the series (4), for $\alpha_{n}=2^{1/n}$, $n=1,2,\ldots,6$. An additional scaling of $(-1)^{n+1}$ is also included so that all solutions are initially positive.
Noteworthy features of the characteristic solutions shown in Fig. 1 include the observation that for $n=1$ the function has no zero in $(0,\infty)$, for $n=2$ there is one zero, for $n=3$ two zeros, and so on. This is reminiscent of Sturm–Liouville theory, but here the roots of two successive characteristic functions do not strictly interlace. (See also [18] for comments about orthogonality, but keep in mind that their definition of an eigenfunction is different, as mentioned below Lemma 1.) Another striking feature is the very flat profile near $t=0$. By repeated differentiation of (3) one can show that if $u(0)=0$, then all derivatives of $u$ vanish at $t=0$. This is suggestive of an essential singularity at the origin, which is consistent with the asymptotic estimate of $u$ as $t\to 0^{+}$ given in [10, Thm. 3]. The asymptotic behaviour as $t\to+\infty$, on the other hand, is given in [10, Thm. 2], namely, $u={\rm O}(e^{-t})$ as $t\to+\infty$. (This follows from (3) under the assumption that $u(\alpha t)\ll u(t)$ for large $t$.)
3 Perturbation analysis of the nonlinear equation
As mentioned in the introduction, the focus of this paper is on the nonlinear problem (2) with initial condition $u(0)=1$. In [7, Thm. 4.1] two families of global solutions have been identified according to the asymptotic limits (A) $u\to 1$ as $t\to\infty$ and (B) $u\to 0$ as $t\to\infty$. Here we restrict ourselves to the cases where these two limits are approached exponentially, i.e., according to (A) $u=1+{\rm O}(e^{-t})$ and (B) $u={\rm O}(e^{-t})$. Algebraic approaches to these limits may also be possible [6], but this will require a different procedure from the one described here.
An obvious solution in class (A) is the constant function $u=1$. Depending on the value of $\alpha$, however, there may be one or more non-constant solutions. When computing these solutions two complications arise, namely, non-linearity and non-uniqueness. A successful computation will therefore require nonlinear iteration starting from a good initial guess. The numerical iteration procedure is described in section 4. Here we present an elementary perturbation analysis for generating good initial guesses, but which may be of interest in its own right.
Because our focus is on solutions in class (A), we define $v(t)=u(t)-1$. Substituting into (2) gives
(7) $v^{\prime}(t)+v(t)=2v(\alpha t)+v^{2}(\alpha t),\quad v(0)=0.$
With the intuition (and support from numerical experimentation) that the characteristic values from the previous section act as bifurcation points in the nonlinear equations (2) and (7), we seek solutions in the neighbourhood of these points by considering $\alpha=\alpha_{n}+\epsilon$, $|\epsilon|\ll 1$. Further, assuming small perturbations about the trivial solution $v(t)=0$, we try the regular expansion $v(t)=\epsilon v_{0}(t)+\epsilon^{2}v_{1}(t)+\cdots$. This gives, to leading order,
(8) $v_{0}^{\prime}(t)+v_{0}(t)=2v_{0}(\alpha_{n}t),\quad v_{0}(0)=0,$
which is the linear equation discussed in section 2.222In an effort to avoid overencumbering notation, we omit burdening $v_{0}$ with an additional subscript, $n$. However, it should be understood here, and throughout, that each $v_{0}$ corresponds to a particular $\alpha=\alpha_{n}$. This value of $n$ should be clear from context. By Lemma 1, the explicit solution to this equation is given by $v_{0}=CE(t;\alpha_{n})$, where $E(t;\alpha_{n})$ is defined in (4). The constant $C$ remains to be determined, however, and this can be done by matching solutions of the linear and nonlinear problems.
To do this, we consider the moments of $v$ as follows. For arbitrary $\alpha$, integrate (7) with respect to $t$ to obtain
[…]
Under the assumption that $v(0)=0$ and $v\to 0$ as $t\to\infty$, the first term vanishes. Making a change of variable $t\mapsto t/\alpha$ on the right, multiplying by $\alpha$, and rearranging gives
[…]
Consider now the particular perturbation solution corresponding to $n=1$, so that $\alpha=2+\epsilon$. Inserting this and $v(t)=\epsilon v_{0}(t)+\epsilon^{2}v_{1}(t)+\cdots$ into (10) and then collecting terms gives, at leading order,
[…]
By plugging in the series solution $v_{0}=C_{1}E(t;2)$ from (4) one finds that the required value of $C_{1}$ is
[…]
where the numerical value has been computed by the formulas (36) in the Appendix. Similar formulas for $C_{n}$, $n>1$, can be derived by considering higher moments of $v$. Their numerical values are given in Table 1 but details of the derivation are deferred to the Appendix.
$n$ $\hskip 16.0ptC_{n}$ $\hskip 13.0pt\|C_{n}E_{n}\|_{2}$
1 11.36911520 1.811978234
2 --809.3665721 9.266935159
3 31551.15567 29.58298033
4 --1099159.137 85.46096526
5 34825078.48 224.4734290
6 --1045480822. 557.2788228
Table 1: The first six scaling coefficients, $C_{1},\ldots,C_{6}$, as given by (33), accurate to all digits shown. Although the $C_{n}$ grow rapidly with $n$, the characteristic functions, $E_{n}(t)=E(t,\alpha_{n})$, are at the same time decreasing in magnitude. Their product, which is the coefficient of $\epsilon$ in the perturbation approximation (13), also increases in magnitude, but not as rapidly as $C_{n}$. Nevertheless, this means that as $\alpha$ approaches the value $1$ with $\epsilon$ fixed, the perturbation approximations become less accurate.
In summary, the approximation to (7) corresponding to $\alpha=2^{1/n}+\epsilon$ is given to first order by
(13) $v(t)\approx C_{n}E(t;2^{1/n})\epsilon,$
where $E(t;2^{1/n})$ is the series solution (4) and the $C_{n}$ values are given in Table 1.
We have not pursued a formal proof of the validity of (13) in any asymptotic sense. (The analysis is rather complicated and perhaps of secondary importance because the principal use of (13) is to provide initial guesses for the numerical method.) Instead, here is a computational check. We substitute (13) into (7), and, expecting the remainder to be ${\rm O}(\epsilon^{2})$, we write
(14) $v^{\prime}(t)+v(t)-2v((2^{1/n}+\epsilon)t)-v^{2}((2^{1/n}+\epsilon)t)=r_{n}(t;\epsilon)\,\epsilon^{2},$
for some coefficient function, $r_{n}$. Every quantity on the left can be explicitly computed from the series (4) and the tabulated values of $C_{n}$. Therefore the values of $r_{n}(t;\epsilon)$ can be computed for each $\epsilon$. By Taylor analysis, the limiting expressions as $\epsilon\to 0$ are found to be
(15) $r_{n}(t;0)=-2C_{n}E^{\prime}(2^{1/n}t;2^{1/n})t-\big{(}C_{n}E(2^{1/n}t;2^{1/n})\big{)}^{2}.$
These functions are shown for a few values of $n$ in Figure 2. Evidently, the $r_{n}(t;0)$ are bounded for all $t\geq 0$, which provides evidence that the approximation (13) satisfies (7) uniformly to ${\rm O}(\epsilon^{2})$. The implied constant grows as $\alpha\to 1$, however, as can be seen by examining the vertical scales in Figure 2. Therefore, the usefulness of (13) as supplier of initial guesses to the numerical method again diminishes accordingly as $n$ increases.
Figure 2: The coefficient functions $r_{n}(t;\epsilon)$ as defined by (14), in the limit as $\epsilon\to 0$. The observation that these functions are bounded on $[0,\infty)$ provides support for the validity of the formulas (13) as uniform approximations to the solutions of (7).
The details of the numerical method are postponed to the next section, but let us now solve (2) (via (7)) for values of $\alpha$ close to the characteristic values. A few such solutions are shown in Figure 3, where the perturbation and numerical approximations are displayed. Evidently, the perturbation approximations provide good qualitative descriptions of the solutions.
We conclude this section with two remarks. Firstly, the $v$ solutions shown in Figure 3 can be distinguished according to whether the first turning point is a local maximum (shown in red) vs a local minimum (shown in blue). This will continue to be a distinguishing factor in the solutions presented below; see for example Figures 5 and 6 in section 5. Secondly, note that the solutions shown in Figure 3 may not be the only ones for the given parameter values. (Additional solutions for $\alpha=\sqrt{2}\pm 0.01$ are shown in subplots J and L of Figure 6.) These additional solutions can be produced by starting from the ones shown in Figure 3 and then using a continuation procedure as described in the next section.
Figure 3: Numerical solutions $v$ of the perturbation equation (7) (solid curves) in comparison to the perturbation approximations (13) (dashed curves). The scaling constants $C_{1}$, $C_{2}$ and $C_{3}$ are given in Table 1. In each frame the value of $\alpha$ is $\alpha_{n}+\epsilon$, with numerical values of these parameters given in the titles and legends.
4 Numerical method
Motivated by the semi-infinite domain, the smoothness and exponential decay of the solutions, and a desire for high accuracy, we solve (7) numerically using a Laguerre spectral collocation method with scaling appropriate to match the $v(t)={\rm O}(e^{-t})$ behaviour for large $t$. To improve efficiency we follow [17] in using a truncated Laguerre method. In particular, we approximate the solution on the first $\lceil M\sqrt{N}\rceil$ points of an $N$-point Laguerre grid, with $M$ determined heuristically based on experiments below. The functional terms, $v(\alpha t)$ and $v^{2}(\alpha t)$, in the equation are implemented with the barycentric resampling approach described in [14]. In short, this approach uses the barycentric interpolation formula to construct a matrix, $P_{\alpha}$, that interpolates the weighted polynomial representation of the solution, $v$, evaluated at the vector of truncated Laguerre points, $\bm{t}$, to the dilated points, $\alpha\bm{t}$, so that $v(\alpha\bm{t})\approx P_{\alpha}v(\bm{t})$. This yields the nonlinear system of equations,
(16) $(D+I-2P_{\alpha})\bm{v}-P_{\alpha}(\bm{v}\circ\bm{v})=0,$
where $\bm{v}$ is the vector of approximate solution values on the truncated Laguerre grid, $D$ and $P_{\alpha}$ are the associated differentiation and barycentric resampling matrices, respectively, and $\circ$ is the element-wise product. To enforce the initial condition, an additional collocation point is included at $t=0$ in the construction of $D$. Because of the boundary condition $v(0)=0$, however, the corresponding row and column of $D$ can subsequently be removed.
We begin by verifying the efficacy of the Laguerre discretisation by solving the linear problem (3) and comparing to the series solution (4). To do so, we omit the nonlinear term from (16) and obtain nontrivial solutions by computing the nullspace of the matrix $D+I-2P_{\alpha}$. Figure 4 shows the error between the spectral collocation and series solutions for various $\alpha_{n}$ as the number of degrees of freedom in the collocation method is increased. The error curves indicate rapid convergence up to a point where the effects of a sub-optimal truncation point degrades the quality of the solution. Rather than trying to find the optimal scaling and discretisation size for each $\alpha$, based on the results in Figure 4, we take $M=6$ and $N=700$ (so that $\lceil M\sqrt{N}\rceil=159$) in our remaining experiments to provide a balance between accuracy and execution time.
Figure 4: Errors in the truncated Laguerre spectral discretisation of (3) for the characteristic solutions at $\alpha=2,2^{1/3},$ and $2^{1/5}$, normalized such that $\|u\|_{2}=1$. Shown are the maximum grid errors as compared to the series solution (4), computed with sufficiently many terms. The series solution is computed in extended precision to mitigate cancellation errors introduced by the large transient growth in its terms for large $n$.
Having verified that the Laguerre discretisation provides accurate solutions to the linear problem (3), we now use it to solve the nonlinear problem (7). We first do so in the vicinity of the characteristic values, $\alpha_{n}=2^{1/n}$. In particular, we use the perturbation approximations derived in section 3 as initial guesses which are then refined using a Newton iteration applied to (16). Some examples have already been shown in Figure 3. For each $n$, these solutions are then used to seed a pseudo-arclength continuation, as described by [9] and implemented in Chebfun [8], to compute solutions over a desired range of $\alpha$.
A MATLAB code that implements the method described here is available at [13]. It can be used to reproduce all figures of this paper.
5 A family of solutions
Using the numerical procedure described in the previous section, we now explore the family of solutions in class (A) as defined in the first paragraph of section 3, i.e., the solutions to (2) that satisfy $u(0)=1$ and $u(t)=1+{\rm O}(e^{-t})$ as $t\to\infty$. Our aim is to catalog for each value of $\alpha$ the number of possible solutions in this class. Recall that for each value of $\alpha$ the constant solution $u=1$ is in this class, so we focus on counting the additional solutions. Excluded in the counting is the solution that satisfies $u(t)={\rm O}(e^{-t})$ as $t\to\infty$, as it is not in class (A). Possible solutions that obey an asymptotic behaviour different from ${\rm O}(e^{-t})$ are excluded in the counting as well (but, as mentioned previously, such solutions are unlikely to be picked up by the numerical method used here).
A bifurcation diagram for solutions in this family is shown in the first frame of Figure 5, and solutions to (2) are shown in the other frames. The bifurcation diagram shows the values of the $L^{2}$-norm (squared) of solutions $v$ to (7) as $\alpha$ is varied. Points at which the curves touch the horizontal axis correspond to $v=0$, therefore representing the constant solution $u=1$. This happens at the critical values of $\alpha$, namely $\alpha_{n}=2^{1/n}$, $n=1,2,\ldots$. In the vicinity of these points, the continuation procedure described in the previous section was used to generate further solutions.
Figure 5: The first frame shows the bifurcation diagram and the remaining frames show solutions to (2) various values of $\alpha$. The black dotted solution decaying to zero is the one in class (B) as defined in the first paragraph of section 3. The multicoloured solutions are in class (A). Note the labeling: the letters at the top of the bifurcation diagram correspond to the various solution plots. Dotted (resp. solid) curves are used to represent solutions whose first turning point is a local maximum (resp. minimum). (An animated version of the figure can be seen at https://youtu.be/aBTvLujm65c.)
Consider first the case $n=1$ (i.e., $\alpha_{1}=2$). Precisely at this value the only solution in class (A) is the constant solution $u=1$ but non-constant solutions emerge if $\alpha$ is perturbed slightly. Such solutions, corresponding to $\alpha=1.9$ and $2.1$, were already displayed in the first frame of Figure 3 (under the translation $u=v+1$). At the end of section 3 we made a remark about distinguishing such solutions according to whether they first dip down or flip up as $t$ is increased. This leads us to the following conventions: Solutions $u$ whose first turning point on the positive $t$-axis is a local maximum will be called plus-solutions and their graphs will be plotted with dotted curves. If a local minimum, these are minus-solutions and plotted with solid curves.
Let us now consider what happens if one decreases the value of $n$ from $1$ down toward 00 (i.e., $\alpha$ increases from the value $2$). Following the first bifurcation curve to the left (dotted blue), one encounters for each value of $\alpha$ a unique plus-solution. An example is the $\alpha=4$ solution shown in the second frame of Figure 5. Increasing the value of $\alpha$ further leads to similar plus-solutions with increasing norm.
Next, consider increasing the value of $n$ from $1$ (i.e., $\alpha$ decreases from $2$). Following the first bifurcation curve to the right (solid blue), one encounters for each value of $\alpha$ a unique minus-solution. An example can be seen in the third frame of Figure 5. When $n$ gets sufficiently close to $2$ ($\alpha$ close to $\sqrt{2}$), however, two additional solutions emerge as represented by the red curve in the bifurcation diagram. Now there are three solutions: the minus-solution of large norm represented by the blue curve as well as the two new solutions, both with relatively small norm.
For more detail about these solutions, consider Figure 6, which zooms in on the interval around $n=2$ in the bifurcation diagram. When $n$ is less than $2$ by a sufficient margin, the only solution in class (A) (other than the constant solution) is a minus-solution such as the one shown with the label H. (This solution is similar to the one computed in [5, Fig. 2] for a slightly different value of $\alpha$.) The solution marked I corresponds to the critical value of $n\approx 1.91$ ($\alpha\approx 1.44$) where a first additional solution is generated. With increasing $n$ this solution bifurcates into two plus-solutions as shown in the subplot labeled J. When $n$ reaches the value $2$, one of these plus-solutions (the one with the smaller norm) momentarily collapses to the constant solution at the bifurcation point (subplot K) and then re-emerges as a minus-solution for larger $n$ (subplot L). Observe that the two solutions of small norm shown in subplots J and L in Figure 6 correspond to the two solutions shown in the middle frame of Figure 3. These are therefore examples of the additional solutions alluded to at the end of section 3.
Figure 6: The bifurcation diagram on the top left is a small section of the one shown in the first frame of Fig. 5. The other frames show the generation of new solutions as the value of $n$ is increased through the value $2$ (i.e., as $\alpha$ is decreased through the value $\sqrt{2}$).
Returning to Figure 5 and continuing to increase the value of $n$, one notices a similar pattern. In each interval $(n-\delta,n)$, with $2\leq n\leq 6$ and $0<\delta\ll 1$, two new solutions are generated. When created, both of these are plus-solutions but one of them switches to a minus-solution once the value of $n$ is exceeded. The magnitude of $\delta$ is about $0.09$ when $n=2$ but then decreases rapidly with $n$. (We conjecture that the inequality $2\leq n\leq 6$ can be replaced by $2\leq n<\infty$. But numerical experimentation in the range $n\geq 6.5$ is complicated by the fact that the perturbation approximation (13) becomes less effective as a starting guess, and therefore non-convergence of the Newton iteration becomes an issue.)
A noteworthy feature of the family of solutions shown in Figure 5 is the fact that the minus-solutions all closely follow the exponentially decaying solution (black dots) for a length of time before turning upwards to approach the value $1$ asymptotically. Other solutions have a tendency to merge as well. The plus-solution that emerges at integer value $n$ catch up to the minus-solution that emerged at integer value $n-1$ as they jointly approach the limiting value $1$.
We conclude with the following puzzling observation. As $\alpha\to 1$, more and more solutions appear. And yet, when $\alpha$ equals $1$ exactly, this multitude of solutions collapses to the unique solution $u=1$. At present we have no good understanding of this nor any means of exploring further because of the difficulty in computing accurate solutions in this regime.
6 Conclusions
The methodology presented here for solving (2) has several ingredients: a Laguerre spectral collocation method for solving the equation numerically, a perturbation method for generating initial guesses for a Newton iteration, and a continuation method for generating further solutions. These techniques are well-established for standard ordinary differential equations, but we believe their combination to solve a nonlinear functional differential equation is new.
The procedure proved to be a reliable way for generating multiple solutions for parameter values $u(0)=1$ and $\alpha>1$, but with $\alpha$ not too close to $1$. Values of $\alpha$ roughly in the range $(1,1.11)$ still provide a challenge and need further investigation.
The methodology should also enable one to consider the following generalization of (2),
(17) $u^{\prime}(t)+au(t)=bu^{m}(\alpha t),$
where $a$ and $b$ are positive constants and $m\geq 2$ an integer. Linearisation about the constant solution $u=(a/b)^{1/m}$ should cast this in a form where the present procedure can be applied with only minor modifications.
Another avenue of investigation is how to solve (2) for other initial conditions. One possibility might be to use the $u(0)=1$ solutions in combination with continuation on the $u_{0}$ parameter to generate solutions corresponding to nearby initial conditions.
7 Acknowledgements
The second author thanks the staff, faculty, and colleagues of the Department of Mathematical Sciences at Stellenbosch University for creating a welcoming and productive environment that made this research possible while on sabbatical leave.
Appendix A Appendix
The details of computing the scaling factors, $C_{n}$, given in Table 1 are as follows. Denote the $j$-th moments of $v$ by
[…]
By multiplying (7) by $t^{j}$ and integrating by parts, it follows that [^3] Here, and below, moments with negative indices are included for ease of notation. They should be taken to be identically zero.
(19) $(\alpha^{j+1}-2)\mu_{j}(v)-j\alpha^{j+1}\mu_{j-1}(v)-\mu_{j}(v^{2})=0,$
where the properties that $v(0)=0$ and $t^{j}v\to 0$ as $t\to\infty$ were used to eliminate the derivative term.
Now, for $n=1,2,\ldots$, consider the perturbation $\alpha=\alpha_{n}+\epsilon$, where $\alpha_{n}=2^{1/n}$, and the expansion $v(t)=\epsilon v_{0}(t)+\epsilon^{2}v_{1}(t)+\cdots$, as in section 3. As before, $v_{0}$ is the solution to (3) with $\alpha=\alpha_{n}$. Substitute this into (19) and use the binomial theorem to expand $(\alpha_{n}+\epsilon)^{j+1}$. Equating the first two powers of $\epsilon$ gives
(20) $(\alpha^{j+1}_{n}-2)\mu_{j}(v_{0})-j\alpha^{j+1}_{n}\mu_{j-1}(v_{0})=0,$
and
(21) $(\alpha^{j+1}_{n}-2)\mu_{j}(v_{1})-j\alpha^{j+1}_{n}\mu_{j-1}(v_{1})+(j+1)\alpha_{n}^{j}\mu_{j}(v_{0})-j(j+1)\alpha_{n}^{j}\mu_{j-1}(v_{0})-\mu_{j}(v_{0}^{2})=0,$
respectively. Using the former expression, the $\mu_{j-1}(v_{0})$ term in the latter can be eliminated, giving
(22) $(\alpha^{j+1}_{n}-2)\mu_{j}(v_{1})-j\alpha^{j+1}_{n}\mu_{j-1}(v_{1})+\frac{2(j+1)}{\alpha}\mu_{j}(v_{0})-\mu_{j}(v_{0}^{2})=0.$
At this stage it is useful to note that, because $v_{0}$ is known—at least, up to the scaling constants $C_{n}$ that we are presently trying to determine—all moments involving $v_{0}$ can be computed numerically; see (36) below. This is not the case for moments of $v_{1}$, however, since $v_{1}$ is unknown, and hence the purpose in the following is to eliminate such moments.
With $j=n-1$, we have $\alpha_{n}^{j+1}=2$ and equations (20) and (22) simplify to
[…]
and
(24) $-2(n-1)\mu_{n-2}(v_{1})+\frac{2n}{\alpha_{n}}\mu_{n-1}(v_{0})-\mu_{n-1}(v_{0}^{2})=0,$
respectively. The fact that the $\mu_{n-1}(v_{1})$ term is eliminated in the latter expression is precisely the reason for considering the $(n-1)$th moment. This will allow us to develop a telescoping recurrence in order to eliminate moments involving the unknown $v_{1}$. To this end, note that (24) can be rearranged as
(25) $\frac{2n}{\alpha_{n}}\mu_{n-1}(v_{0})=\mu_{n-1}(v_{0}^{2})+2(n-1)\mu_{n-2}(v_{1}),$
which we return to momentarily.
Now consider $j<n-1$. It follows from (20) and (23) that
(26) $\mu_{j}(v_{0})=0,\qquad j<n-1.$
This causes the third term on the left of (22) to vanish, which gives the recurrence
(27) $\mu_{j}(v_{1})=\frac{1}{\alpha_{{n}}^{j+1}-2}\left(\mu_{j}(v_{0}^{2})+j\alpha_{{n}}^{j+1}\mu_{j-1}(v_{1})\right),\qquad j<n-1.$
Returning now to (25) and applying (27) with $j=n-2$ gives
(28) $\frac{2n}{\alpha_{n}}\mu_{n-1}(v_{0})=\mu_{n-1}(v_{0}^{2})+\frac{2(n-1)}{\alpha^{n-1}_{n}-2}\left(\mu_{n-2}(v_{0}^{2})+(n-2)\alpha_{n}^{n-1}\mu_{n-3}(v_{1})\right).$
If $n=2$ the term on the right involving $v_{1}$ is eliminated and, with $\alpha_{2}=2^{1/2}$, one then gets
(29) $2\sqrt{2}\,\mu_{1}(v_{0})=\mu_{1}(v_{0}^{2})-(2+\sqrt{2})\,\mu_{0}(v_{0}^{2}).$
If $n>2$ then further substitutions of (27) into (28) are made until the resulting series likewise terminates with the vanishing of the quantity involving $v_{1}$ and only moments of $v_{0}$ and $v_{0}^{2}$ remain. The general formula involves the following linear combination of the $\mu_{j}(v_{0}^{2})$ on the right
(30) $\frac{2n}{\alpha_{n}}\mu_{n-1}(v_{0})=\sum_{j=0}^{n-1}w_{j}\mu_{j}(v_{0}^{2}),\qquad n=1,2,\ldots,$
where the $w_{j}$ are generated in the reverse direction by $w_{n-1}=1$, and
(31) $w_{j-1}=\frac{j\alpha^{j+1}_{n}}{\alpha^{j}_{n}-2}w_{j},\quad j=n-1,\ldots,1.$
Letting $v_{0}(t)=C_{n}E(t;\alpha_{n})$ as in (4), it follows that (30) becomes
(32) $\frac{2n}{\alpha_{n}}C_{n}\mu_{n-1}(E_{n})=C_{n}^{2}\sum_{j=0}^{n-1}w_{j}\mu_{j}(E^{2}_{n}),$
where we have denoted $E(t;\alpha_{n})$ by $E_{n}$ for brevity. Finally, solving for $C_{n}$ gives
(33) $C_{n}=\frac{2n\mu_{n-1}(E_{n})}{\alpha_{n}\displaystyle\sum_{j=0}^{n-1}w_{j}\mu_{j}(E^{2}_{n})},\quad n=1,2,\ldots.$
For $n=1$ ($\alpha_{1}=2$) one gets
(34) $C_{1}=\frac{\mu_{0}(E_{1})}{\mu_{0}(E_{1}^{2})},$
a formula already given in (12). For $n=2$ ($\alpha_{2}=\sqrt{2}$) the formula is
(35) $C_{2}=\frac{2\sqrt{2}\,\mu_{1}(E_{2})}{\mu_{1}(E^{2}_{2})-(2+\sqrt{2})\mu_{0}(E_{2}^{2})},$
which also follows directly from (29). For larger values of $n$ the formulas are not as clean as the surds do not cancel elegantly.
Explicit series formulas for the moments in (33)–(35) can be obtained by integrating the series (4) term-by-term. In particular,
(36) $\mu_{j}(E_{n})=j!\sum_{k=0}^{\infty}\frac{b_{k}}{\alpha_{n}^{(j+1)k}}\qquad\text{and}\qquad\mu_{j}(E^{2}_{n})=j!\sum_{k=0}^{\infty}\sum_{\ell=0}^{\infty}\frac{b_{k}b_{\ell}}{(\alpha_{n}^{k}+\alpha_{n}^{\ell})^{j+1}},$
where
(37) $b_{k}=\left\{\begin{array}[]{cl}1,&k=0,\\ \displaystyle\frac{2^{k}}{(1-\alpha_{n})\cdots(1-\alpha_{n}^{k})},&k>0.\end{array}\right.$
Although the coefficients in these series have exponential asymptotic decay, they have alternating signs and large transient growth, making them numerically unstable, particularly when $n$ is large. Extended precision is therefore required to compute the moments accurately from these formulas. Alternatively, one can approximate the moments with numerical integration, but care is still required in evaluating (4).